### v2

@@ 2 Discussion
Large Language Models hold substantial promise for the medical domain, particularly in augmenting digital workflows and improving efficiency in healthcare delivery. The ability to integrate external tools and functionality with LLMs paves the way for innovative applications and can overcome limitations of LLMs, such as the potential to hallucinate.
@@
We have demonstrated how LLMs can provide a unique interface between healthcare professionals and clinical predictive models, such as risk scores. Our approach is scalable and requires no additional training, although approaches that improve with use could be yet more powerful. Additionally, we aim to mitigate the problem of LLM hallucination and ensure that actionable advice is anchored in approved clinical resources.
@@
While LLMs have general capabilities, they can lack domain-specific knowledge. This has led to the development of medical-focused LLMs, either by training new LLMs from scratch Luo2022 ; taylor2022galactica ; Yang2022large or by adapting existing general-purpose LLMs Singhal2023 . While we showed using such specialist LLMs is not required, they are easy to incorporate due to the modularity of the approach and doing so could further enhance the functionality of LLM-based interfaces.
@@
As AI in medicine continues to advance, further research into LLMs and their potential applications is not only beneficial but vital. LLMs could help to alleviate the data burden that is contributing to clinician burnout, as well as streamline patient management processes. Furthermore, studies have demonstrated high usability of LLMs, even with limited experience Skjuve2023user , which is critical for successful clinical deployment. While we believe this paper represents an important first step, we are only scratching the surface of the potential of LLMs in healthcare. Ultimately, this line of work may significantly change the digital health landscape, enhancing the capabilities of clinicians and the quality of patient care.
@@ 3 Code availability
Code for the LLM-based interfaces can be accessed at https://github.com/pauliusrauba/LLMs_interface.
@@ 4 Data availability
The risk predictor for diabetes was trained on a cohort from the UK Biobank resource. Data from UK Biobank is accessible through a request process (https://www.ukbiobank.ac.uk/enable-your-research/register). The authors had no special access or privileges when accessing the data.
@@
    A.1 Large Language Models
    By default, LLMs will provide responses in the form of text based on the provided context, such as a prompt or conversation history. To construct interfaces for digital tools using LLMs, we instead view the LLM as an agent that can interact with an environment to solve tasks. Formally, at each step $t\in T$, the agent receives observation $o_{t}\in O$ from the environment and subsequently takes action $a_{t}\in A$ according to policy $\pi(a_{t}|h_{t})$, where $h_{t}=(o_{0},a_{0},\dots,o_{t-1},a_{t-1},o_{t})$ is the history. To enable the agent to both reason and use external tools, we use the ReAct framework yao2022react which decomposes the action space as $\hat{A}=A\cup L$, where $a\in A$ are actions using specific tools and an action $a\in L$ represents not using an external tool but instead allows the model to reason over the history about what action to take next.
    While frameworks such as Toolformer schick2023toolformer and WebGPT nakano2021webgpt trained LLMs to use basic tools, such as calculators, calendars, and search engines, via self-supervised fine-tuning and fine-tuning using behavior cloning and reinforcement learning, respectively. Following ReAct yao2022react , we used in-context learning dong2023survey in the form of prompting, providing the LLM with sufficient information about possible actions and using the underlying reasoning capabilities of the LLM. This removes the need for further training of the LLM, which might be challenging in the medical domain without suitable examples. This also enabled us to use multiple tools. We built user interfaces using StreamLit streamlit .
    A.2 External Tools
    This section describes the external tools and sources of information made available to the LLM in our two examples. In the subsequent section, we provide the system and tool prompts.
    CVD Risk Prediction. In our first example, we show how LLMs can incorporate existing tools and information for CVD risk prediction (Fig. 3a). We provided the LLM access to QRisk3 Hippisley-Cox2017CVD , a risk prediction tool that assesses the likelihood of developing CVD within 10 years. QRISK3 is currently the recommended risk score for CVD in the United Kingdom and is widely used to guide decisions about who might benefit from interventions NICE_CVD . We enabled the LLM to use the risk score either using the provided data or, if requested by the user, to modify a variable and assess the impact of such a change on the patient’s risk. Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD .
    Diabetes Risk Prediction. To demonstrate an application for diabetes risk prediction using a machine learning model (Fig. 3b), we equipped the LLM with the following tools. For the diabetes risk score, we trained a predictive model using AutoPrognosis 2.0 Imrie2023AutoPrognosis on a cohort of individuals from UK Biobank Sudlow2015 with no known history of diabetes at baseline. The primary outcome was diagnosis of diabetes within a 10-year horizon. We selected 20 features based on their predictive value Imrie2023AutoPrognosis , namely HbA1c, blood glucose, sex, age, waist size, hip size, weight, BMI, waist-height ratio, waist-hip ratio, gamma-glutamyl transferase, cystatin-c, C-reactive protein, atrial fibrillation, HDL, alanine transaminase, triglycerides, uric acid, sex hormone-binding globulin, use of anti-hypertensives.
    Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 ; asan2020artificial . A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking . To enable this, we enabled the LLM to use explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions, both in general and for the specific patient. In particular, we used SHAP lundberg2017unified to interpret model predictions.
    Figure S.1: Illustration of the process by which the LLM uses external tools. The LLM is provided with the history of the interaction, including the current request. Using an iterative reasoning process, the LLM decides which, if any, tools are required and with what input (“Thought”). The LLM then uses the external tool (“Action”) and receives the output (“Observation”). Finally, the LLM decides to answer the question (“Thought”), providing a response to the user.

### v3

@@ 3 Discussion
Large Language Models hold substantial promise for the medical domain, particularly in augmenting digital workflows and improving efficiency in healthcare delivery. The ability to integrate external tools and functionality with LLMs paves the way for innovative applications and can overcome limitations of LLMs, such as hallucinations. Doing so offers a potential transformation for how clinicians interact with digital tools and could help overcome the challenges of deploying clinical AI models.
@@
We have demonstrated how LLMs can provide a unique interface between healthcare professionals and clinical predictive models, such as risk scores. In particular, we developed a novel risk score for incidence of CVD using automated machine learning and developed LLM-based interfaces for our model and QRisk3 Hippisley-Cox2017CVD , the current recommended risk score in the UK for CVD. Our approach is scalable and does not require any additional training of the language model, although approaches that improve with use could be yet more powerful. Additionally, we aim to mitigate the problem of hallucination by ensuring that actionable advice is anchored in approved clinical resources, contrasting several previous applications of LLMs in medicine that focused exclusively on the knowledge and information learned by LLMs.
Numerous considerations exist before, during, and after assessing an individual’s prognosis (Table 2). Currently, clinicians must access these tools via fixed user interfaces or application programming interfaces (APIs), with existing interfaces typically only calculating risk. Through an LLM-based interface, practitioners can obtain substantial additional information about the risk score, its development and methodology, the prediction issued, and related medical guidelines, in a manner that specifically addresses their needs or questions without providing superfluous information.
In this paper, we have focused on clinicians interacting with digital tools. However, there are numerous stakeholders in healthcare in addition to clinicians, such as patients, regulators, and administrators, each with different goals and requirements Imrie2023 . Our framework and approach could be applied to improve digital health interfaces for these alternate stakeholders. While this could have additional challenges, there are potentially even more substantial benefits for such individuals, given the differences in requirements, knowledge, and familiarity with digital health technology, among other factors.
@@
While LLMs have general capabilities, they can lack domain-specific knowledge. This has led to the development of medical-focused LLMs, either by training new LLMs from scratch Luo2022 ; taylor2022galactica ; Yang2022large or by adapting existing general-purpose LLMs Singhal2023 . While we showed using such specialist LLMs is not required, they could be readily incorporated due to the modularity of our approach. This could further enhance the functionality of LLM-based interfaces.
@@
As AI in medicine continues to advance, further research into LLMs and their potential applications in healthcare could provide significant benefits. For example, LLMs could help to alleviate the data burden that is contributing to clinician burnout, as well as streamline patient management processes. Furthermore, studies have demonstrated high usability of LLMs, even with limited experience Skjuve2023user , which is critical for successful clinical deployment. While we believe this paper represents an important first step, we are only scratching the surface of the potential of LLMs in healthcare. Ultimately, this line of work may significantly change the digital health landscape, enhancing the capabilities of clinicians and the quality of patient care.
4 Methods
4.1 Data source and study population
We developed and validated our prognostic model using data from UK Biobank Sudlow2015 , a large prospective cohort of individuals from the UK. Participants in UK Biobank were enrolled between 2006 and 2010 and aged between 40 and 70 at the time of recruitment. We extracted a cohort of individuals with no known history of CVD at baseline. We excluded individuals who had been diagnosed with chronic kidney disease or type 1 diabetes and individuals who were being treated with statins. This resulted in a cohort of 407,605 individuals (Fig. S.2).
4.2 Clinical Predictors and Outcomes
We began by considering a number of predictors previously validated as CVD risk factors. A subset of features were selected based on their predictive power and feature importance. We included 20 features in our model, namely sex, age, atrial fibrillation, steroid treatment, type 2 diabetes, hypertension treatment, family history of CVD, systolic blood pressure, smoking status (never-smoker, ex-smoker, light smoker, moderate smoker, heavy smoker, alkaline phosphatase, apolipoprotein a, apolipoprotein b, cystatin c, C-reactive protein, HbA1c, IGF-1, lipoprotein a, triglycerides, urea, and overall health rating, which was self-reported by participants as excellent, good, fair or poor.
The primary outcome was incidence of CVD within a 10-year horizon. This was defined using the following ICD-10 codes: myocardial infarction (I21, I22), stroke (I63, I64), angina pectoris (I20), or transient cerebral ischaemic attacks (G45). There were 17,600 CVD events within a 10-year horizon in the UK Biobank cohort.
4.3 Model Derivation
We trained a prognostic model using AutoPrognosis 2.0, an open-source automated machine learning software package Imrie2023AutoPrognosis . AutoPrognosis has been validated in several applications in medicine, for example to determine eligibility for lung cancer screening callender2023assessing . AutoPrognosis was used to optimize pipelines consisting of a variable preprocessing step followed by model selection and training. The optimized pipelines were subsequently combined in an ensemble using a weighted combination. Individual pipelines and the final ensemble were trained to maximize discrimination, measured using the area under the receiver operating curve. Predictions from the final ensemble were calibrated in a piecewise linear manner using the observed risk calculated using a Kaplan-Meier estimator kaplan1958nonparametric .
4.4 Statistical Analysis
Model discrimination was assessed using the concordance index (C-index) Uno2011 and calibration via calibration curves and the ratio between the expected and observed number of events. We also assessed Brier scores brier1950verification . Calibration curves were constructed by first dividing the population into 10 risk deciles based on each individual’s predicted 10-year risk and then comparing the mean estimated risk to the observed risk in each decile. Kaplan–Meier estimators kaplan1958nonparametric were used to calculate observed risk. Additionally, we conducted decision curve analysis, assessing the net benefit for a range of risk thresholds. In addition to the existing risk scores, when assessing net benefit, we also compared our model with baseline strategies that assume all patients develop CVD (All) or none do (None). Results are based on 5-fold cross-validation. We used multiple imputation with chained equations (MICE) vanBuuren2011 to impute missing values. We generated five imputed datasets and combined results using Rubin’s rules rubin2018multiple .
4.5 Model Comparisons
We compared our model with three existing risk scores, namely Framingham score DAgostino2008 , SCORE2 SCORE2 , and QRisk3 Hippisley-Cox2017CVD . All three models assess the likelihood of developing CVD within a 10 year horizon.
The Framingham score DAgostino2008 was recommended by the 2010 American College of Cardiology/American Heart Association (ACC/AHA) guidelines Greenland2010 and is based on eight risk factors: sex, age, systolic blood pressure, treatment for hypertension, smoking status, history of diabetes, total cholesterol, and high-density lipoprotein (HDL) cholesterol.
SCORE2 SCORE2 was developed using data from 45 cohorts in 13 countries across Europe, including the United Kingdom, and has been included in the European Society of Cardiology guidelines. SCORE2 employs seven risk factors: sex, age, systolic blood pressure, smoking status, history of diabetes, total cholesterol, and high-density lipoprotein (HDL) cholesterol.
QRisk3 Hippisley-Cox2017CVD was developed on data from 981 practices in England comprising almost 8 million patients and is currently the recommended risk score for CVD in England NICE_CVD . QRisk3 uses 22 variables, including demographics (such as sex, age), lifestyle factors (e.g. smoking), medical and family history (e.g. history of diabetes, family history of CVD), medical examination (e.g. blood pressure), and blood tests (total cholesterol/HDL cholesterol ratio).
4.6 LLM-based interface
Here, we describe the implementation details for the LLM-based systems demonstrated above.
4.6.1 LLM framework
By default, LLMs will provide responses in the form of text based on the provided context, such as a prompt or conversation history. To construct interfaces for digital tools using LLMs, we instead viewed the LLM as an agent that can interact with an environment to solve tasks. Formally, at each step $t\in T$, the agent receives observation $o_{t}\in O$ from the environment and subsequently takes action $a_{t}\in A$ according to policy $\pi(a_{t}|h_{t})$, where $h_{t}=(o_{0},a_{0},\dots,o_{t-1},a_{t-1},o_{t})$ is the history. To enable the agent to both reason and use external tools, we used the ReAct framework yao2022react which decomposes the action space as $\hat{A}=A\cup L$, where $a\in A$ are actions using specific tools and an action $a\in L$ represents not using an external tool but instead allows the model to reason over the history about what action to take next.
Frameworks such as Toolformer schick2023toolformer and WebGPT nakano2021webgpt trained LLMs to use basic tools, such as calculators, calendars, and search engines, via self-supervised fine-tuning and fine-tuning using behavior cloning and reinforcement learning, respectively. In contrast, following ReAct yao2022react , we used in-context learning dong2023survey in the form of prompting, providing the LLM with sufficient information about possible actions and using the underlying reasoning capabilities of the LLM. This removes the need for further training of the LLM, which might be challenging in the medical domain without suitable examples. This also enabled us to use multiple tools. We built user interfaces using StreamLit streamlit .
4.6.2 External Tools
This section describes the external tools and sources of information made available to the LLM in our examples.
LLM-based interface for AutoPrognosis. To demonstrate an LLM-based interface, we equipped the LLM with the CVD risk prediction model developed in this work on the UK Biobank cohort using AutoPrognosis 2.0 Imrie2023AutoPrognosis . Building model trust is a crucial step for prognostic models, in particular for models that are not inherently interpretable Rajpurkar2022 ; asan2020artificial . A recent study found that medical decision-makers had a strong preference for interactive explanations and, in particular, for these interactions to take the form of natural language dialogues lakkaraju2022rethinking . To enable this, we enabled the LLM to use explainable AI (XAI) methods Imrie2023 on the underlying model, allowing users to investigate the rationale for predictions, both in general and for the specific patient. In particular, we used SHAP lundberg2017unified to interpret model predictions. We additionally provided the LLM with information about the variables used by the risk score, the underlying methodology and how the model was constructed, and details regarding the cohort used to develop the model. Example interactions are shown in Fig. 6 and Fig. S.3.
LLM-based interface for QRisk3. As a second example, we showed how LLMs can incorporate existing tools and information for CVD risk prediction. We provided the LLM access to QRisk3 Hippisley-Cox2017CVD and enabled the LLM to use the risk score either using the provided data or, if requested by the user, to modify a variable and assess the impact of such a change on the patient’s risk. Additionally, we provided the LLM with access to the academic paper describing QRisk3 Hippisley-Cox2017CVD and the National Institute for Health and Care Excellence (NICE) clinical guidelines for CVD NICE_CVD . An example interaction is shown in Fig. 7.
@@ 5 Code availability
Code for the LLM-based interfaces can be accessed at https://github.com/pauliusrauba/LLMs_interface. AutoPrognosis is an open-source package available on GitHub at https://github.com/vanderschaarlab/AutoPrognosis and on PyPI at https://pypi.org/project/autoprognosis/.
@@ 6 Data availability
This research has been conducted using the UK Biobank resource under application number 105160. Data from UK Biobank is accessible through a request process (https://www.ukbiobank.ac.uk/enable-your-research/register). The authors had no special access or privileges when accessing the data.